OpenStudy (anonymous):

In an electric field of strength 0.50 N/C, a test charge experiences a force of 2.50 × 10^-4 N in the same direction as the field. What is the magnitude of the test charge? A. -5.00 × 10^-4 C B. -2.50 × 10^-4 C C. +2.50 × 10^-4 C D. +5.00 × 10^-4 C

OpenStudy (michele_laino):

the requested charge (magnitude) is given by the subsequent formula: \[\large Q = \frac{F}{E} = \frac{{2.5 \times {{10}^{ - 4}}}}{{0.5}}\]

OpenStudy (anonymous):

ok! so we get 5E-4? so our answer is choice D? :O

OpenStudy (michele_laino):

that's right! Since vectors E and Q have the same orientation 

OpenStudy (anonymous):

ahh okie!! yay!! thank you!:)

OpenStudy (michele_laino):

vectors E and F

OpenStudy (michele_laino):

:)